The Warren Watch Company sells watches for $30, fixed costs are $150,000, and variable costs are $11 per watch.

  1. What is the firm's gain or loss at sales of 6,000 watches? Enter loss (if any) as negative value. Round your answer to the nearest cent.

    What is the firm's gain or loss at sales of 16,000 watches? Enter loss (if any) as negative value. Round your answer to the nearest cent.
  2. What is the break-even point (unit sales)? Round your answer to the nearest whole number.
    units
  3. What would happen to the break-even point if the selling price was raised to $35?
    The result is that the break-even point is lower? The result is that the break-even point is higher? The result is that the break-even point remains unchanged?
  4. What would happen to the break-even point if the selling price was raised to $35 but variable costs rose to $26 a unit? Round your answer to the nearest whole number.
    The result is that the break-even point increases? The result is that the break-even point decreases? The result is that the break-even point remains unchanged?
Answer #1

Gain = Sales – variable costs – Fixed costs

a.Gain = (30-11)*6,000 – 150,000

= -$36,000

i.e. loss

b.Gain = 16000*19 – 150000

= $154,000

i.e. gain

c.Break even unit sales = Fixed costs/(Selling price per watch – variable cost per watch)

= 150,000/19

= 7,894.74 watches

c. The result is that the break-even point is lower. Since contribution margin will increase

d.The result is that the break even point is higher. since contribution margin will decrease.